# Fundamental physical constants set the observability and operation of phase and other transitions and increase entropy

### v3

@@ Abstract
Approaching the problem of understanding fundamental physical constants (FPCs) started with discussing the role these constants play in high-energy nuclear physics and astrophysics. Condensed matter physics was relatively unexplored in this regard. More recently, it was realised that FPCs set lower or upper bounds on key condensed matter properties. In this Perspective, we discuss a wider role played by FPCs in condensed matter physics: at given environmental conditions, FPCs set the observability and operation of entire physical effects and phenomena. We discuss second-order phase transitions including structural and superconducting transitions and first-order transitions including melting and boiling. We also discuss metastable states, transitions between them, chemical reactions and their products. An interesting byproduct of this discussion is that the order of magnitude of the transition temperature can be calculated from FPCs. We show that the new states emerging as a result of various transitions increase the phase space and entropy. Were FPCs to take different values, these transitions would become inoperative at our environmental conditions and the new states due to these transitions would not emerge. Generalising this insight to other levels of structure in the Universe, this suggests that the current values of FPCs, by enabling various transitions and reactions which give rise to new states, act to increase the phase space and entropy.
@@ 1 Introduction
Fundamental physical constants (FPCs) determine physical processes at all energy and length scales of our experiments and theories. These different length scales are illustrated in Figure 1. The values of FPCs are often listed in textbooks and reviews (see, e.g., Refs. Uzan (2003, 2011); Tiesinga et al. (2021); Ashcroft and Mermin (1976)), and their recommended values are updated and maintained in the National Institute for Standards and Technology database NIST .
@@
Together with our fundamental theories, FPCs are the key to understanding the present state of our Universe and its contents as well as its past and future Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al. (2022); Cahn (1996); Hogan (2000); Adams (2019); Uzan (2003). The “barcodes of ultimate reality” Barrow (2003), FPCs also set our Universe apart from those worlds where these constants might be different.
@@
Understanding the values and origin of FPCs is one of the grandest challenges in modern science Barrow and Webb (2006). The problem of understanding fundamental constants is not that it involves very hard technical calculations but that we do not know anything more fundamental Weinberg (1983). We do not know what kind of theoretical tools we need to address the problem. It is probably this enormity of the problem that explains why there seems to be no sustained programme of research in this area, scheduled meetings and so on, as compared to more developed fields of science.
@@
A starting point to understand FPCs is to ask what would happen were FPCs to take values different from those currently observed Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al. (2022); Cahn (1996); Hogan (2000); Adams (2019)? This is done on the basis of our physical models. The bottom in Figure 1 involves length scales of particle physics. According to our theories, observed particles involve a finely-tuned balance between the masses of down and up quarks: larger up-quark mass gives the neutron world without protons and hence no atoms made from nuclei and electrons around them, whereas larger down-quark mass gives the proton world without neutrons where light atoms can form only but not heavy atoms. Our world with heavy atoms and electrons endowing complex chemistry would not exist with a few per cent fractional change in the mass difference of the two quarks Hogan (2009, 2000).
@@
The top in Figure 1 involves cosmological scales and a balance between the gravitational constant and the cosmological constant. This balance gives enough time for the Universe to expand and not to contract too quickly and at the same time maintain sufficient gravitational clamping so that galaxies, stars and other structures can form Barrow (2003); Adams (2019).
@@
For the rest of length scales in Figure 1, our theories imply that there is an intricate network of inter-related constraints on FPCs that need to be satisfied to give many structures in the observable Universe. This includes formation and evolution of different astronomical objects and in particular stars, synthesis of heavy elements in stars including carbon and oxygen and other elements which can then form molecular structures essential to life Barrow (2003); Adams (2019). The strength of these constraints, the degree of tuning between FPCs, depends on how well developed our physical models are. Ref. Adams (2019) discusses existing uncertainties involved in our models and how these uncertainties affect constraints imposed on FPC. It is important to note that FPCs involved in these processes are often the same and include the Planck constant, speed of light in vacuum, electron mass and charge and the proton-to-electron mass ratio (this is different from the constants operating at the very top and bottom scales mentioned earlier - these play role at their length and energy scales only). This means that the parameter space of these FPCs can be better constrained and understood.
@@
It was realised that FPCs set upper or lower bounds on several key condensed matter properties. Examples include bounds on viscosity Trachenko and Brazhkin (2020), diffusion and spin dynamics Luciuk et al. (2017); Sommer et al. (2011); Bardon et al. (2014); Trotzky et al. (2015); Enss and Thywissen (2018), electron and heat transport Hartnoll (2015); Mousatov and Hartnoll (2020); Behnia and Kapitulnik (2019); Nussinov and Chakrabarty (2022), thermal conductivity in insulators Trachenko et al. (2021) and conductors Brazhkin (2023), the speed of sound Trachenko et al. (2020), elastic moduli including those in lower-dimensional systems Brazhkin (2023); Trachenko (2023a), velocity gradients that can be set up in cells using biochemical energy Trachenko (2023b) and superconducting temperature Trachenko et al. . This has deepened our understanding of these properties and their physical models.
@@
We show that the new states emerging from various transitions increase the phase space and entropy. We also show that were FPCs to take different values, these transitions would become inoperative at our environmental conditions, and the new states due to these transitions would not emerge. Generalising this insight to other levels of structure in the Universe, this suggests that the current values of FPCs, by enabling various transitions and reactions which give rise to new states, act to increase the phase space and entropy. This invites an analogy with an evolutionary process where entropy increase selects for FPCs.
@@ 2 Why filling the condensed matter gap is important
Second, recall that it is the uncertainty and insufficient development of physical models related to higher-energy processes that limits our understanding of FPCs Adams (2019). On the other hand, condensed matter physics involves a larger variety of ways to test condensed matter theories. Many of these tests are more readily available than those in high energy physics. As a result, theories and concepts in condensed matter physics have been well verified and cross-checked using a great variety of experiments and simulations.
@@ 3.1 Fundamental and environmental parameters
We start with separating two types of parameters. The first type are FPCs. The second type are the environmental parameters such as temperature and pressure at typical planetary conditions where we do our observations. These parameters are not fundamental but are due to historical accidents Weinberg (1993) such as the distance between the planet and its star which affects the planet’s temperature. In several examples in this section, we consider a certain range of environmental parameters ($\Delta T$,$\Delta P$) around, for example, temperature (102-103 K) and pressure range at typical planetary conditions in the Solar system and ask how the observability and operation of condensed matter phenomena change as a result of different FPCs.
We will be using and discussing the Planck constant $\hbar$, electron mass $m_{e}$ and charge $e$ and the proton mass $m_{p}$. We will also be discussing two dimensionless ratios, the proton-to-electron mass ratio $\beta$
@@
$\alpha$ and $\beta$ are important in key processes at different length and energy scales in Figure 1, including the stability of particles, nuclear reactions, formation and evolution of stars, synthesis of heavy nuclei and so on. These processes impose constraints on the allowed range of $\alpha$ and $\beta$ (as well as on the strong nuclear constant) Barrow (2003); Adams (2019).
@@
$a_{\rm B}\approx 0.5$ and is on the short side of interatomic separations in condensed phases in the typical range 1-3 Å. $E_{\rm R}\approx$ 13.6 eV and is on the large side of cohesive energies in condensed phases in the range 1-10 eV.
@@ 3.2.1 Enropy increase due to a phase transition
Our first case study are second-order phase structural transitions in solids. Let’s consider common silica SiO2 and its low-temperature phase, quartz. Silica has several polymorphs at different pressure and temperature, separated by second-order phase transitions. As a model example, let’s the high-temperature phase of silica, $\beta$-cristobalite, emerging at the phase transition Hammonds et al. (1996); Dove (1997).
The solid entropy in the classical high-temperature regime $T\gg\hbar\omega_{\rm D}$, where $\omega_{\rm D}$ is Debye frequency, is Landau and Lifshitz (1970):
@@
As expected for high-temperature phases on general grounds Landau and Lifshitz (1970), cristobalite has a more symmetric and open structure compared to quartz. As a result, cristobalite has a significantly larger number of low-frequency phonons, rigid-unit modes, that propagate without the distortion of rigid SiO4 tetrahedra Hammonds et al. (1996); Dove (1997). This gives smaller $\bar{\omega}$ in cristobalite and larger entropy in Eq. (5).
@@
Frequency softening above the phase transition can be seen more generally in terms of the order parameter $\eta$. Considering that phonon frequencies are renormalised by anharmonicity related to a phase transition, the renormalised frequencies $\tilde{\omega}_{i}$ are Dove (1997, 2003):
@@
The entropy increase can also take place below $T_{c}$ if the high-temperature phase becomes a long-lived metastable state below $T_{c}$. In our model quartz-cristobalite example, the phase transition is a reconstructive phase transition with large energy barriers separating the two phases. As a result, cristobalite is a long-lived metastable phase which can be cooled back to room temperature or any other temperature below the phase transition Hammonds et al. (1996); Dove (1997). According to Eq. (5), there is no net entropy change due to heating quartz from temperature $T_{1}$ to the transition temperature $T_{2}$ and cooling back the metastable cristobalite phase: the entropy first increases by $3N\ln\frac{T_{2}}{T_{1}}$ and then decreases by the same amount (we do not consider the small effect of phonon frequency softening with temperature due to usual anharmonicity). Before proceeding further, we add a note about the entropy and other thermodynamic functions of non-equilibrium metastable systems. These functions are defined in the commonly used sense: at times shorter than the relaxation time of the non-equilibrium system, this system can be conditionally considered in an equilibrium state where observed properties are described by equilibrium thermodynamics Landau and Lifshitz (1970). This is discussed in Section 3.5 in more detail.
We find both quartz and cristobalite at our environmental conditions, hence let’s consider the entropy of the quartz-cristobalite mixture at room or any other temperature below $T_{c}$ where both phases co-exist, $S_{Q-C}$. This mixture consists of any proportion of the two phases with the total number of atoms $N$. We compare $S_{Q-C}$ to the entropy of the initial quartz phase, $S_{Q}$, with the same number of atoms $N$. $S_{Q-C}$ is larger than $S_{Q}$ because the entropy of the cristobalite phase is larger as discussed earlier. Hence,
@@
The increased phase space and entropy in the high-temperature phase become “frozen-in” in the metastable state below $T_{c}$.
@@
We now observe that if we consider different values of FPCs, the phase transition would not operate. The upper bound on the maximal phonon frequency $\omega_{m}$ (“cage rattling” frequency) can be written in terms of fundamental constants Trachenko et al. . This bound follows from combining Eq. (4) and the ratio of $\hbar\omega_{m}$ and cohesive energy $E$ Trachenko and Brazhkin (2020)
@@
$\omega_{m}$ has the upper bound, $\omega_{m}^{u}$, for $m=m_{p}$ as Trachenko et al.
@@
$\omega_{m}^{u}$ interestingly contains the proton-to-electron mass ratio $\beta$ (1). As mentioned earlier, this ratio is of particular importance for different processes and reactions involving different length scales in Figure 1 Barrow (2003); Adams (2019).
Lets consider what would happen to a phase transition were FPCs to take different values and in particular consider variations of FPCs which increase $\omega_{m}$ and $\omega_{m}^{u}$ in Eqs. (12)-(13). Note that this question is well-posed because it is possible to vary $\omega_{m}$ and $\omega_{m}^{u}$ by varying $\hbar$, $m_{e}$, $e$ and $m_{p}$ while keeping $\beta$ and $\alpha$ in Eqs. (1) and (2) unchanged and hence maintain all preceding essential processes (e.g. those involved in stellar formation and evolution, synthesis of heavy elements, stability of ordered structures and so on) intact Barrow (2003); Adams (2019).
Large $\omega_{m}$ implies large $\omega_{i}$ and $\omega_{\rm D}$ (recall that $\omega_{m}$ is close to $\omega_{\rm D}$) and $T\ll\hbar\omega_{\rm D}$ in the temperature range considered. This, in turn, implies that the high-temperature classical result (5) does not apply because the system is now in the quantum regime. In this regime, instead of Eq. (5), the entropy is Landau and Lifshitz (1970)
@@
The observability of phase transitions can also be discussed on the basis of $T_{c}$. In phase transition theories, including the Ising model and the anharmonic phonon theory (equivalent to Landau theory), the order of magnitude of $T_{c}$ is set by the energy of interaction between particles Dove (1997). This energy is on the order of the phonon energy $\hbar\omega_{\rm D}$ or $\hbar\omega_{m}$. Using Eq. (12), this gives
@@
$A^{\frac{1}{2}}$ varies across the periodic table in the range of about 1 to 15, with an average value of 8. Using this value and $\omega_{m}^{u}=3680$ K in Eq. (13), Eq. (16) gives $T_{c}\approx$ 500 K. This is close to $T_{c}$ calculated from typical parameters of the anharmonic phonon model Dove (1993) and is of the right order of magnitude of $T_{c}$ in the range $10^{2}-10^{3}$ K typically observed.
We see that the order of magnitude of $T_{c}$ (15) is set by FPCs. Were different values of FPCs to give $T_{c}$ on the order of, for example, $10^{5}$ K or higher, phase transitions would be unobservable at our environmental conditions. This is consistent with our earlier argument involving small entropy in Eq. (14) due to large $\omega_{\rm D}$.

### v4

@@ Abstract
Approaching the problem of understanding fundamental physical constants (FPCs) started with discussing the role these constants play in high-energy nuclear physics and astrophysics. Condensed matter physics was relatively unexplored in this regard. More recently, it was realised that FPCs set lower or upper bounds on key condensed matter properties. In this Perspective, we discuss a wider role played by FPCs in condensed matter physics: at given environmental conditions, FPCs set the observability and operation of entire physical effects and phenomena. We discuss second-order phase transitions including structural and superconducting transitions and first-order transitions including melting and boiling. We also discuss metastable states, transitions between them, chemical reactions and their products. An interesting byproduct of this discussion is that the order of magnitude of the transition temperature can be calculated from FPCs. We show that the new states emerging as a result of various transitions increase the phase space and entropy. Were FPCs to take different values, these transitions would become inoperative at our environmental conditions and the new states due to these transitions would not emerge. This suggests that the current values of FPCs, by enabling various transitions and reactions which give rise to new states, promote entropy increase. We propose that this is a general effect that applies to different levels of structure in the Universe.
@@ 1 Introduction
Fundamental physical constants (FPCs) determine physical processes at all energy and length scales of our experiments and theories. These different length scales are illustrated in Figure 1. The values of FPCs are often listed in textbooks and reviews (see, e.g., Refs. [1, 2, 3, 4]), and their recommended values are updated and maintained in the National Institute for Standards and Technology database [5].
@@
Together with our fundamental theories, FPCs are the key to understanding the present state of our Universe and its contents as well as its past and future [6, 7, 8, 9, 10, 11, 12, 13, 1]. The “barcodes of ultimate reality” [6], FPCs also set our Universe apart from those worlds where these constants might be different.
@@
Understanding the values and origin of FPCs is one of the grandest challenges in modern science [14]. The problem of understanding fundamental constants is not that it involves very hard technical calculations but that we do not know anything more fundamental [15]. We do not know what kind of theoretical tools we need to address the problem. It is probably this enormity of the problem that explains why there seems to be no sustained programme of research in this area, scheduled meetings and so on, as compared to more developed fields of science.
@@
A starting point to understand FPCs is to ask what would happen were FPCs to take values different from those currently observed [6, 7, 8, 9, 10, 11, 12, 13]? This is done on the basis of our physical models. The bottom in Figure 1 involves length scales of particle physics. According to our theories, observed particles involve a finely-tuned balance between the masses of down and up quarks: larger up-quark mass gives the neutron world without protons and hence no atoms made from nuclei and electrons around them, whereas larger down-quark mass gives the proton world without neutrons where light atoms can form only but not heavy atoms. Our world with heavy atoms and electrons endowing complex chemistry would not exist with a few per cent fractional change in the mass difference of the two quarks [16, 12].
@@
The top in Figure 1 involves cosmological scales and a balance between the gravitational constant and the cosmological constant. This balance gives enough time for the Universe to expand and not to contract too quickly and at the same time maintain sufficient gravitational clamping so that galaxies, stars and other structures can form [6, 13].
@@
For the rest of length scales in Figure 1, our theories imply that there is an intricate network of inter-related constraints on FPCs that need to be satisfied to give many structures in the observable Universe. This includes formation and evolution of different astronomical objects and in particular stars, synthesis of heavy elements in stars including carbon and oxygen and other elements which can then form molecular structures essential to life [6, 13]. The strength of these constraints, the degree of tuning between FPCs, depends on how well developed our physical models are. Ref. [13] discusses existing uncertainties involved in our models and how these uncertainties affect constraints imposed on FPC. It is important to note that FPCs involved in these processes are often the same and include the Planck constant, speed of light in vacuum, electron mass and charge and the proton-to-electron mass ratio (this is different from the constants operating at the very top and bottom scales mentioned earlier - these play role at their length and energy scales only). This means that the parameter space of these FPCs can be better constrained and understood.
@@
It was realised that FPCs set upper or lower bounds on several key condensed matter properties. Examples include bounds on viscosity [17], diffusion and spin dynamics [18, 19, 20, 21, 22], electron and heat transport [23, 24, 25, 26], thermal conductivity in insulators [27] and conductors [28], the speed of sound [29], elastic moduli including those in lower-dimensional systems [28, 30], velocity gradients that can be set up in cells using biochemical energy [31] and superconducting temperature [32]. This has deepened our understanding of these properties and their physical models.
@@
We show that the new states emerging from various transitions increase the phase space and entropy. We also show that were FPCs to take different values, these transitions would become inoperative at our environmental conditions, and the new statess would not emerge. This suggests that the current values of FPCs, by enabling various transitions and reactions which produce the new states, result in the increase of the phase space and entropy. We propose that this is a general effect that applies to other levels of structure in the Universe. This discussion invites an analogy with an evolutionary process where entropy increase selects for FPCs.
@@ 2 Why filling the condensed matter gap is important
Second, recall that it is the uncertainty and insufficient development of physical models related to higher-energy processes that limits our understanding of FPCs [13]. On the other hand, condensed matter physics involves a larger variety of ways to test condensed matter theories. Many of these tests are more readily available than those in high energy physics. As a result, theories and concepts in condensed matter physics have been well verified and cross-checked using a great variety of experiments and simulations.
@@ 3.1 Fundamental and environmental parameters
We start with separating two types of parameters. The first type are FPCs. The second type are the environmental parameters such as temperature $T$ and pressure $P$ at typical planetary conditions where we do our observations. These parameters are not fundamental but are due to historical accidents [33] such as the distance between the planet and its star which affects the planet’s temperature. In the Solar system, for example, $T$ is in the range 102-103 K. In this section, we will consider how FPCs affect the observability and operation of condensed matter phenomena such as phase transitions and other reactions at our environmental conditions.
We will be discussing the Planck constant $\hbar$, electron mass $m_{e}$ and charge $e$ and the proton mass $m_{p}$. We will also be discussing two dimensionless ratios, the proton-to-electron mass ratio $\beta$
@@
$\alpha$ and $\beta$ are important in key processes at different length and energy scales in Figure 1, including the stability of particles, nuclear reactions, formation and evolution of stars, synthesis of heavy nuclei and so on. These processes impose constraints on the allowed range of $\alpha$ and $\beta$ (as well as on the strong coupling constant) [6, 13].
@@
$a_{\rm B}\approx 0.5$ and is on the shorter side of interatomic separations in condensed phases in the typical range 1-3 Å. $E_{\rm R}\approx$ 13.6 eV and is on the larger side of cohesive energies in condensed phases in the range 1-10 eV.
@@ 3.2.1 Enropy increase due to a phase transition
Our first case study are second-order phase structural transitions in solids. Let’s consider common silica SiO2 and its low-temperature phase, quartz. Silica has several polymorphs at different pressure and temperature, separated by second-order phase transitions. As a model example, let’s the high-temperature phase of silica, $\beta$-cristobalite, emerging at the phase transition [34, 35].
The solid entropy in the classical high-temperature regime $T\gg\hbar\omega_{\rm D}$, where $\omega_{\rm D}$ is Debye frequency, is [36]:
@@
As expected for high-temperature phases on general grounds [36], cristobalite has a more symmetric and open structure compared to quartz. As a result, cristobalite has a significantly larger number of low-frequency phonons, rigid-unit modes, that propagate without the distortion of rigid SiO4 tetrahedra [34, 35]. This gives smaller $\bar{\omega}$ in cristobalite and larger entropy in Eq. (5).
@@
Frequency softening above the phase transition can be seen more generally in terms of the order parameter $\eta$. Considering that phonon frequencies are renormalised by anharmonicity related to a phase transition, the renormalised frequencies $\tilde{\omega}_{i}$ are [35, 37]:
@@
The entropy increase can also take place below $T_{c}$ if the high-temperature phase is a long-lived metastable state below $T_{c}$. In our quartz-cristobalite example, the phase transition is a reconstructive transition with large energy barriers separating the two phases. As a result, cristobalite is a long-lived metastable phase which can be cooled back to room or any other temperature below $T_{c}$ [34, 35]. We note here that the entropy and other thermodynamic functions of non-equilibrium metastable systems are defined in the commonly used sense: at times shorter than the relaxation time of the non-equilibrium system, this system can be conditionally considered in an equilibrium state where observed properties are described by equilibrium thermodynamics [36]. This is discussed in Section 3.5 in more detail.
We find both quartz and cristobalite at our environmental conditions, hence let’s consider the entropy of the quartz-cristobalite mixture at any temperature below $T_{c}$ where both phases co-exist, $S_{Q-C}$. This mixture consists of any proportion of the two phases with the total number of atoms $N$. We compare $S_{Q-C}$ to the entropy of the initial quartz phase, $S_{Q}$, with the same number of atoms $N$. $S_{Q-C}$ is larger than $S_{Q}$ because the entropy of cristobalite is larger as discussed earlier. Hence,
@@
The increased phase space and entropy above $T_{c}$ become “frozen-in” in the metastable state below $T_{c}$.
@@
We now observe that if we consider different values of FPCs, the phase transition would not operate. The upper bound on the maximal phonon frequency $\omega_{m}$ (“cage rattling” frequency) can be written in terms of fundamental constants [32]. This bound follows from combining Eq. (4) and the ratio of $\hbar\omega_{m}$ and cohesive energy $E$ [17]
@@
$\omega_{m}$ has the upper bound, $\omega_{m}^{u}$, for $m=m_{p}$ as [32]
@@
$\omega_{m}^{u}$ interestingly contains the proton-to-electron mass ratio $\beta$ (1). As mentioned earlier, this ratio is of particular importance for different processes and reactions involving different length scales in Figure 1 [6, 13].
Let’s consider what would happen to a phase transition were FPCs to take different values. Let’s consider variations of FPCs which increase $\omega_{m}$ and $\omega_{m}^{u}$ in Eqs. (12)-(13). Note that this question is well-posed because it is possible to vary $\omega_{m}$ and $\omega_{m}^{u}$ by varying $\hbar$, $m_{e}$, $e$ and $m_{p}$ while keeping $\beta$ and $\alpha$ in Eqs. (1) and (2) unchanged and hence maintain all preceding essential processes (e.g. stellar formation and evolution, nuclear synthesis, stability of ordered structures and so on) intact [6, 13].
Large $\omega_{m}$ implies large $\omega_{i}$ and $\omega_{\rm D}$ (recall that $\omega_{m}$ is close to $\omega_{\rm D}$) and $T\ll\hbar\omega_{\rm D}$ in the temperature range considered. This, in turn, implies that the high-temperature classical result (5) does not apply because the system is now in the quantum regime. In this regime, instead of Eq. (5), the entropy is [36]
@@
The observability of phase transitions can also be discussed on the basis of $T_{c}$. Phase transition theories, including the Ising model and the anharmonic phonon theory (equivalent to Landau theory), expectedly predict that the order of magnitude of $T_{c}$ is set by the energy of interaction between particles [35]. This energy is close to the phonon energy $\hbar\omega_{\rm D}$ or $\hbar\omega_{m}$. Using Eq. (12),
@@
$T_{c}$ can be estimated by recalling that $A^{\frac{1}{2}}$ varies across the periodic table in the range of about 1-15, with an average of 8. Using this value and $\omega_{m}^{u}=3680$ K (13), Eq. (16) gives $T_{c}\approx 500$ K. This is close to $T_{c}$ derived using typical parameters of the anharmonic phonon model and Landau theory [38] and is of the right order of magnitude of $T_{c}$ typically observed in the range $10^{2}$-$10^{3}$ K.
We see that the order of magnitude of $T_{c}$ (15) is set by FPCs. Were different values of FPCs to give $\omega_{m}$ and $T_{c}$ in Eq. (15) on the order of, for example, $10^{5}$ K or higher, phase transitions would be unobservable at our environmental conditions. This is consistent with our earlier result: large $\omega_{\rm D}$ gives small entropy in Eq. (14) and no phase transition.